# IF-RewardBench: Benchmarking Judge Models for Instruction-Following Evaluation

### 05 Mar 2026 v1

@@ 1 Introduction
We conduct a comprehensive evaluation of 21 popular judge models on IF-RewardBench, including state-of-the-art dedicated reward models and general LLMs. Our results highlight a substantial capability gap: even the leading proprietary LLM, Gemini-3-Pro, achieves only a moderate Kendall correlation of 0.609 in ranking responses, significantly falling behind the human performance of 0.755. Top-tier open-source models, such as GLM-4.6 and Deepseek-V3.2, remain below or near 0.4, while all dedicated reward models fail to exceed 0.2. In-depth analysis reveals several critical insights: (1) Although constraint-level scoring outperforms overall pairwise comparison, the limited capacity for detecting constraint violation remains a primary performance bottleneck. (2) Incorporating system instructions or conversation history exacerbates evaluation difficulty for overall pairwise comparison. (3) Situation and Style constraints that involve subjectivity prove harder to verify than objective ones. (4) Judge performance degrades with increased constraint composition complexity, number of constraints within instructions, and response quality, underscoring the growing challenge of evaluating advanced models in complex application scenarios. Notably, IF-RewardBench presents greater difficulty than existing benchmarks and demonstrates a significantly stronger positive correlation with the downstream BoN sampling performance of judge models. These findings establish IF-RewardBench as a valuable resource for advancing instruction-following evaluation.
@@ 3.3 Dataset Stastics
Table 2: Statistics of IF-RewardBench, including the number of instructions (#Inst.), the average dialog turns (#Turn), the number of constraints (#Cons.) per instruction, the average number of responses (#Resp. and preference relations (#Rela.) per preference graph.
The final IF-RewardBench contains 842 instructions across three types, as shown in Table 2. For each instruction, we collect $m=8$ responses from a single LLM, filtering out those unrelated to any preference relations. On average, each preference graph contains 7.14 responses and 10.86 preference relations, highlighting its structural complexity. Among these responses, 74.6% of constraints are annotated as "followed". Our annotation process achieves a Cohen’s Kappa Cohen (1960) of 0.67 in initial instruction-following annotation and 0.87 in cross-validation, indicating almost perfect agreement and high data quality. Unlike existing benchmarks that mainly consider code-verifiable constraints and their And composition, IF-RewardBench covers a more diverse spectrum of constraints, as shown in Figure 3. We also provide a length difference analysis in Appendix F.
4 Experiments
@@ 4.1 Evaluation Setup
For constraint assessment, we compute the F1-scores for both positive and negative classes in each example to assess verification capability, capturing both error detection and the avoidance of over-criticism. The Kendall ($\tau_{b}$) correlation coefficient between the quality scores derived from Equation 1 and the golden truth preference relations in each preference graph is used to assess ranking capability. For overall assessment, we also adapt the Kendall ($\tau_{b}$) correlation coefficient as the evaluation metric for ranking capability.
@@
We evaluate a total of 21 popular judge models, including state-of-the-art dedicated reward models and general LLMs, as listed in Appendix G.1. General LLMs are applicable to both constraint and overall assessment via elaborate prompting, while dedicated reward models are only suitable for overall assessment. For general LLMs in overall assessment, we adapt the widely used pairwise comparison paradigm. In this paradigm, we follow RRM Guo et al. (2025b) and compare all possible response pairs to calculate the ELO Elo (1978) scores for ranking multiple responses. More implementation details are in Appendix G.2.
@@ Appendix A Details of Constraint Taxonomy
To comprehensively evaluate the capability of judge models in various instruction-following scenarios, we first propose a hierarchical constraint taxonomy that delineates both constraint categories and their composition types. For the former, we merge the categories from several previous works Qin et al. (2024); Jiang et al. (2024); Qin et al. (2025a); Li et al. (2025b) to establish a comprehensive taxonomy comprising 7 primary constraint categories and 54 secondary categories. For the latter, we adapt the taxonomy from ComplexBench Wen et al. (2024) to define 4 constraint composition types. The detailed description and corresponding example of each primary constraint category and constraint composition type are presented in Table 7 and 8, respectively.
@@ B.2 Instruction Synthesis
Given the difficulty of simulating realistic user behaviors in multi-turn interaction scenarios, we only leverage LLMs to synthesize complex instructions in the single-turn interaction scenario. During the instruction synthesis step, we randomly sample 2 to 5 secondary constraint categories from the taxonomy described in Appendix A and prompt Deepseek-R1 Guo et al. (2025a) to inject several corresponding constraints into the seed instructions, thereby elevating their complexity. Unlike previous works that primarily focus on independent and atomic constraints Ren et al. (2025); Peng et al. (2025a), we further instruct Deepseek-R1 to combine multiple constraints using 4 composition types: Single, And, Chain, Selection, and the nested structure of the above composition types. The detailed definition of these composition types is provided in Appendix A. In this way, our instruction synthesis framework can capture the multifaceted instruction types in real-world scenarios more comprehensively. Finally, we employ Deepseek-R1 to validate the generated instructions, ensuring they are reasonable, unambiguous, and logically consistent.
@@ Appendix C List for Prompt Templates
This section lists all applied prompt templates throughout this work, including the prompt for scoring the quality and difficulty of user instructions in Table 16 and 17, the prompt for instruction decomposition in Table 18, the prompt for constraint assessment in Table 19, and the prompt for pairwise comparison in overall assessment in Table
@@ E.1 Annotation Guidelines
For instruction-following annotation, we provide annotators with an instruction, the corresponding constraint checklist, and the model response. Human annotators are instructed to judge whether the model response follows each constraint in the checklist and provide a brief justification. The detailed guideline is shown in Table 15.
@@ G.1 Evaluated Judge Models
As shown in Table 11, we evaluate a total of 21 popular judge models on IF-RewardBench, including representative general LLMs, fine-tuned discriminative reward models, and generative reward models.
@@ 2. If it’s negative, analyze the reasons for it:
Interaction Type Source
Single-Turn Interaction InfoBench Qin et al. (2024), FollowBench Jiang et al. (2024), ComplexBench Wen et al. (2024), Inverse-IFEval Zhang et al. (2025a), IFBench Pyatkin et al. (2025), CFBench Zhang et al. (2025b), Meeseeks Zhao et al. (2025)
Multi-Turn Interaction Multi-IF He et al. (2024), MultiChallenge Deshpande et al. (2025), StructFlowBench Li et al. (2025a)
System-Prompt Steerability RoleMRC Lu et al. (2025), SysBench Qin et al. (2025a), IHEval Zhang et al. (2025d), AgentIF Qi et al. (2025)
Table 9: The source instruction-following benchmarks for the instructions in IF-RewardBench.
@@
Qwen-2.5-72B-Instruct Yang et al. (2024) 72B Alibaba
Qwen-2.5-7B-Instruct Yang et al. (2024) 7B Alibaba
@@
Model BoN BoN LLMBar RB2 IFBench IF-RewardBench IF-RewardBench
(CA) (OA) (CA) (OA)
Gemini-3-Flash 0.851 0.854 0.887 0.794 0.802 0.572 0.513
GPT-5-mini 0.838 0.857 0.853 0.713 0.757 0.519 0.456
DeepSeek-V3.2 0.820 0.835 0.881 0.681 0.678 0.395 0.288
GLM-4.6 0.817 0.837 0.868 0.513 0.617 0.422 0.270
GLM-4.5-Air 0.815 0.826 0.824 0.506 0.640 0.308 0.148
QwQ-32B 0.824 0.821 0.846 0.531 0.637 0.356 0.183
Qwen-3-32B 0.811 0.819 0.821 0.381 0.644 0.285 0.129
Qwen-3-8B 0.808 0.816 0.787 0.400 0.601 0.230 0.097
Llama-3.3-70B-Instruct 0.810 0.809 0.831 0.381 0.624 0.238 0.054
Llama-3.1-8B-Instruct 0.811 0.796 0.439 0.331 0.577 0.089 0.000
Qwen-2.5-72B-Instruct 0.808 0.803 0.705 0.419 0.572 0.181 0.048
Qwen-2.5-7B-Instruct 0.807 0.815 0.611 0.294 0.552 0.094 0.041
Skywork-Reward-V2-Llama-3.1-8B - 0.817 0.881 0.644 0.628 - 0.133
Llama-3.1-70B-Instruct-RM-RB2 - 0.804 0.815 0.388 0.599 - 0.124
InternLM2-20B-Reward - 0.798 0.718 0.344 0.631 - 0.040

### 16 Apr 2026 v2

@@ 1 Introduction
We conduct a comprehensive evaluation of 22 popular judge models on IF-RewardBench, including state-of-the-art dedicated reward models and general LLMs. Our results highlight a substantial capability gap: even the leading proprietary LLM, Gemini-3-Pro, achieves only a moderate Kendall correlation of 0.609 in ranking responses, significantly falling behind the human performance of 0.755. Top-tier open-source models, such as GLM-4.6 and Deepseek-V3.2, remain below or near 0.4, while all dedicated reward models fail to exceed 0.2. In-depth analysis reveals several critical insights: (1) Although constraint-level scoring outperforms overall pairwise comparison, the limited capacity for detecting constraint violation remains a primary performance bottleneck. (2) Incorporating system instructions or conversation history exacerbates evaluation difficulty for overall pairwise comparison. (3) Situation and Style constraints that involve subjectivity prove harder to verify than objective ones. (4) Judge performance degrades with increased constraint composition complexity, number of constraints within instructions, and response quality, underscoring the growing challenge of evaluating advanced models in complex application scenarios. Notably, IF-RewardBench presents greater difficulty than existing benchmarks and demonstrates a significantly stronger positive correlation with the downstream BoN sampling performance of judge models. These findings establish IF-RewardBench as a valuable resource for advancing instruction-following evaluation.
@@ 3.2.2 Preference Graph Curation
Table 2: Statistics of IF-RewardBench, including the number of instructions (#Inst.), the average dialog turns (#Turn), the number of constraints (#Cons.) per instruction, the average number of responses (#Resp.) and preference relations (#Rela.) per preference graph.
@@
3.3 Dataset Stastics
The final IF-RewardBench contains 842 instructions across three types, as shown in Table 2. For each instruction, we collect $m=8$ responses from a single LLM, filtering out those responses unrelated to any preference relations. On average, each preference graph contains 7.14 responses and 10.86 preference relations, highlighting its structural complexity. Among these responses, 74.6% of constraints are annotated as "followed". Our annotation process achieves a Cohen’s Kappa coefficient Cohen (1960) of 0.67 in initial instruction-following annotation and 0.87 in cross-validation, indicating almost perfect agreement and high data quality. Unlike existing benchmarks that mainly consider code-verifiable constraints and their And composition, IF-RewardBench covers a more diverse spectrum of constraints categories and their composition types, as shown in Figure 3. We also conduct a length difference analysis in Appendix F and confirm that the preference relations are not confounded by length bias Zheng et al. (2023).
4 Experiments
@@ 4.1 Evaluation Setup
For constraint assessment, we compute the F1-scores for both positive and negative classes in each example to assess verification capability, capturing both error detection and the avoidance of over-criticism. The Kendall ($\tau_{b}$) correlation coefficient between the quality scores derived from Equation 1 and the golden truth preference relations in each preference graph is used to assess ranking capability. For overall assessment, we also calculate the Kendall ($\tau_{b}$) correlation coefficient as the evaluation metric for ranking capability.
@@
We evaluate a total of 22 popular judge models, including state-of-the-art dedicated reward models and general LLMs, as listed in Appendix G.1. General LLMs are applicable to both constraint and overall assessment via elaborate prompting, while dedicated reward models are only suitable for overall assessment. For general LLMs in overall assessment, we adapt the widely used pairwise comparison paradigm. In this paradigm, we follow RRM Guo et al. (2025b) and compare all possible response pairs to calculate the ELO Elo (1978) scores for ranking multiple responses. More implementation details are in Appendix G.2.
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Science Foundation for Distinguished Young Scholars (with No. 62125604) and the Natural Science Foundation of China (No. 62536008). We would also like to thank Zhipu AI for sponsoring the computational resources and annotation costs in this work.
@@ Appendix A Details of Constraint Taxonomy
To comprehensively evaluate the capability of judge models in various instruction-following scenarios, we first propose a hierarchical constraint taxonomy that delineates both constraint categories and their composition types. For the former, we merge the categories from several previous works Qin et al. (2024); Jiang et al. (2024); Qin et al. (2025a); Li et al. (2025b) to establish a comprehensive taxonomy comprising 7 primary constraint categories and 54 secondary categories. For the latter, we adapt the taxonomy from ComplexBench Wen et al. (2024) to define 4 constraint composition types. The detailed description and corresponding example of each primary constraint category and constraint composition type are presented in Table 8 and 9, respectively.
@@ B.2 Instruction Synthesis
Given the difficulty of simulating realistic user behaviors in multi-turn interaction scenarios, we only leverage LLMs to synthesize complex instructions in the single-turn interaction scenario. The seeds are collected from an online LLM-based chat service platform that serves more than a million users daily. We utilize LLMs to automatically score the difficulty of these seed instructions and retain those with lower difficulty to facilitate constraint synthesis. During the instruction synthesis step, we randomly sample 2 to 5 secondary constraint categories from the taxonomy described in Appendix A and prompt Deepseek-R1 Guo et al. (2025a) to inject several corresponding constraints into the seed instructions, thereby elevating their complexity. Unlike previous works that primarily focus on independent and atomic constraints Ren et al. (2025); Peng et al. (2025a), we further instruct Deepseek-R1 to combine multiple constraints using 4 composition types: Single, And, Chain, Selection, and the nested structure of the above composition types. The detailed definition of these composition types is provided in Appendix A. In this way, our instruction synthesis framework can capture the multifaceted instruction types in real-world scenarios more comprehensively. Finally, we employ Deepseek-R1 to validate the generated instructions, ensuring they are reasonable, unambiguous, and logically consistent.
@@ Appendix C List for Prompt Templates
This section lists all applied prompt templates throughout this work, including the prompt for scoring the quality and difficulty of user instructions in Table 17 and 18, the prompt for instruction decomposition in Table 19, the prompt for constraint assessment in Table 20, and the prompt for pairwise comparison in overall assessment in Table
@@ E.1 Annotation Guidelines
For checklist revision, we provide annotators with an instruction and the corresponding constraint checklist. Annotators are instructed to check the correctness of the checklist and correct any errors. The detailed guideline is shown in Table 15. For instruction-following annotation, we provide annotators with an instruction, the corresponding constraint checklist, and the model response. Annotators are instructed to judge whether the model response follows each constraint in the checklist and provide a brief justification. The detailed guideline is shown in Table 16.
@@ G.1 Evaluated Judge Models
As shown in Table 12, we evaluate a total of 22 popular judge models on IF-RewardBench, including representative general LLMs, fine-tuned discriminative reward models, and generative reward models.
@@ G.2 Implementation Details of Judge Models
Model BoN BoN LLMBar RB2 IFBench IF-RewardBench IF-RewardBench
(CA) (OA) (CA) (OA)
Gemini-3-Flash 0.851 0.854 0.887 0.794 0.802 0.572 0.513
GPT-5-mini 0.838 0.857 0.853 0.713 0.757 0.519 0.456
DeepSeek-V3.2 0.820 0.835 0.881 0.681 0.678 0.395 0.288
GLM-4.6 0.817 0.837 0.868 0.513 0.617 0.422 0.270
GLM-4.5-Air 0.815 0.826 0.824 0.506 0.640 0.308 0.148
QwQ-32B 0.824 0.821 0.846 0.531 0.637 0.356 0.183
Qwen-3-32B 0.811 0.819 0.821 0.381 0.644 0.285 0.129
Qwen-3-8B 0.808 0.816 0.787 0.400 0.601 0.230 0.097
Llama-3.3-70B-Instruct 0.810 0.809 0.831 0.381 0.624 0.238 0.054
Llama-3.1-8B-Instruct 0.811 0.796 0.439 0.331 0.577 0.089 0.000
Qwen-2.5-72B-Instruct 0.808 0.803 0.705 0.419 0.572 0.181 0.048
Qwen-2.5-7B-Instruct 0.807 0.815 0.611 0.294 0.552 0.094 0.041
Skywork-Reward-V2-Llama-3.1-8B - 0.817 0.881 0.644 0.628 - 0.133
Llama-3.1-70B-Instruct-RM-RB2 - 0.804 0.815 0.388 0.599 - 0.124
InternLM2-20B-Reward - 0.798 0.718 0.344 0.631 - 0.040
@@ 2. If it’s negative, analyze the reasons for it:
Instruction Type Source # Instruction
Single-Turn Interaction InfoBench Qin et al. (2024) 21
FollowBench Jiang et al. (2024) 15
ComplexBench Wen et al. (2024) 38
Inverse-IFEval Zhang et al. (2025a) 25
IFBench Pyatkin et al. (2025) 90
CFBench Zhang et al. (2025b) 25
Meeseeks Zhao et al. (2025) 15
Real Application Scenarios 64
LLM Synthesis 100
Multi-Turn Interaction Multi-IF He et al. (2024) 78
MultiChallenge Deshpande et al. (2025) 22
StructFlowBench Li et al. (2025a) 12
Real Application Scenarios 90
System-Prompt Steerability RoleMRC Lu et al. (2025) 40
SysBench Qin et al. (2025a) 100
IHEval Zhang et al. (2025d) 50
AgentIF Qi et al. (2025) 7
Real Application Scenarios 50
Table 10: The sources of the instructions in IF-RewardBench.
@@
Qwen-2.5-72B-Instruct Yang et al. (2024a) 72B Alibaba
Qwen-2.5-7B-Instruct Yang et al. (2024a) 7B Alibaba
@@
Below, an instruction that may contain the system prompt and conversation history, and a constraint checklist for the instruction will be provided. The checklist includes all the constraints the model must follow when generating a response. Your task is to verify whether the checklist accurately and comprehensively covers all constraints, and to modify it if any errors are found.
Constraint Definition:
Constraints refer to all instructional directives from the system prompt, conversation history, and user instructions that apply to the final-round response, excluding auxiliary information such as background knowledge, reference texts, and in-context examples. They include, but are not limited to, descriptions of the basic tasks to be completed and specific requirements for output content, style, and format.
Task Details:
Step 1: Error Identification
Evaluate the accuracy of the checklist. Select "Incorrect" if any of the following conditions are met:
1. Vague Constraints: The checklist includes ambiguous constraints, such as "satisfy all constraints mentioned above" or "your response must be correct".
2. Omissions: The checklist misses mandatory constraints present in the system prompt, conversation history, or user instruction.
3. Expired Constraints: The checklist includes outdated constraints from the history or system prompt that no longer apply to the current round.
4. Fabricated Constraints: The checklist contains constraints not presented in the system prompt, history, or user instruction.
5. Incorrect Paraphrasing: The checklist inaccurately interprets a constraint (e.g., the original instruction states "list the key points in an ordered list," but the checklist incorrectly paraphrases it as "list the key points in an unordered list").
Step 2: Checklist Modification
If you select "Incorrect" in Step 1, modify the checklist to make it correct. You must follow these principles:
1. Source Integrity: All checklist items must originate from the provided information.
2. Completeness and Accuracy: Do not omit necessary constraints, duplicate existing ones, or fabricate new ones.
3. Atomicity: Each constraint must be atomic yet possess complete semantics, ensuring it can be fully understood without referencing other constraints.
{examples}
[System Prompt]
{system_prompt}
[Conversation History]
{history}
[Final Round User Instruction]
{user_prompt}
[Constraint Checklist]
{checklist}
Your judgment of the checklist quality: {option} A. Correct  B. Incorrect
Your modified checklist (Required if B is selected): {modified-checklist}
Table 15: Human annotation guideline for checklist revision. The red part is the information provided to the annotators, and the blue part is content that requires the annotators to make annotations.
@@
[The Start of Final Round User Instruction]
{user_prompt}
[The End of Final Round User Instruction]